OpenStudy (thats_kyy):

A pencil at a stationery store costs $1, and a pen costs $1.50. Stefan spent $7 at the store. He bought a total of 6 items. Which system of equations can be used to find the number of pencils (x) and pens (y) he bought? x + 1.5y = 7 x + y = 6 1.5x + y = 7 x = 6y x + 2y = 7 6x = 1.5y 2x + y = 7 6x = 1.5y

OpenStudy (thats_kyy):

@geerky42

OpenStudy (thats_kyy):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

x = # of pencils y = # of pens he bought 6 items (pens & pencils) so what equation can we form?

OpenStudy (thats_kyy):

6x + 6=y @jim_thompson5910

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

(number of pencils) + (number of pens) = total number of items

jimthompson5910 (jim_thompson5910):

use that template

OpenStudy (thats_kyy):

so 6+6 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

x = # of pencils y = # of pens

jimthompson5910 (jim_thompson5910):

6 is the total number of items bought

OpenStudy (thats_kyy):

6x + 6=y

jimthompson5910 (jim_thompson5910):

no, it should be x+y = 6

jimthompson5910 (jim_thompson5910):

you might be thinking of something else?

OpenStudy (thats_kyy):

sorry i can end o be slow at times

OpenStudy (thats_kyy):

tend*

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (thats_kyy):

okay So the equation will be x + y =6

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

that's one of the equations in the system

OpenStudy (thats_kyy):

ok i think i have the other one so im guessing my answer will be A @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (thats_kyy):

Thank You Your A Huge Help!!

jimthompson5910 (jim_thompson5910):

yw